# Intrinsic Geometry of the Stock Market from Graph Ricci Flow

### v1

@@ Abstract
We use the discrete Ollivier-Ricci graph curvature with Ricci flow to examine the intrinsic geometry of financial markets through the empirical correlation graph of the NASDAQ 100 index. Our main result is the development of a technique to perform surgery on the neckpinch singularities that form during the Ricci flow of the empirical graph, using the behavior and the lower bound of curvature of the fully connected graph as a starting point. We construct an algorithm that uses the curvature generated by intrinsic geometric flow of the graph to detect hidden hierarchies, community behavior, and clustering in financial markets despite the underlying challenges posed by a highly connected geometry.
@@
Edge-centric approaches involving graph curvature are increasingly emerging as the method of choice to analyze complex graphs. Combinatorial Forman-Ricci[19],[14],[20] curvature and probabilistic Ollivier-Ricci curvature[8],[9],[10] (ORC) are key developments that have permitted the adaptation of differential geometry methods to discrete structures such as graphs. The correlation matrix, being a fully connected graph, already provides an important indicator of the geometry of the system. Graph curvature has been evoked in the context of the stock market but in these works the network was again replaced by the MST[21],[22],[23],[24]. In this letter, we use the Ollivier-Ricci curvature to construct and analyze the weighted network generated by the stock market as a weighted complete graph. Our main result is to propose a discrete Ricci flow algorithm with surgery based on curvature along neckpinch singularities which allows us to analyze the structure and intrinsic geometry of fully connected graphs with positive curvature, such as the stock market. Detailed comparisons with state-of-the art techniques such as MST and PCA are the subject of a longer paper[25].
@@
From this definition, we see that if the nodes $i$ and $j$ belong to the same community, they tend to share neighbors. Therefore the cost of transporting mass is smaller and the Wasserstein distance is no greater than $d(i,j)$. Thus intra-community edges have a positive Ollivier-Ricci curvature. On the other hand, if nodes $i$ and $j$ are from different communities, they tend to have fewer neighbors in common. Therefore, transporting mass from $i$ to $j$ necessarily involves passing through the edge $ij$ and the Wasserstein distance is larger than $d(i,j)$. Therefore the curvature of edge $ij$ is negative and the link is hyperbolic. On a flat graph, the Wasserstein and $d(i,j)$ distances are identical and the curvature is zero. If all the links in a graph have zero curvature, then the graph is called Ricci-flat. The regular grid is an example of a Ricci-flat graph.
@@
In Fig. 1 we present our construction of the NASDAQ-100 graph with empirical edge weights, constructed from real-life data. We recovered the values of stock prices from Yahoo! Finance[32] using the components of the NASDAQ-100 Index from [33]. We calculate the returns and average over a period of five years from 1/1/2020-20/11/2024. Based on the edge weights, we obtain a histogram of Ricci curvatures[27],[34]. The histogram is shifted downward from the complete graph at 0.505, though being all positive retain the spherical aspect of this graph. This is reflected in the downward shift of the distances from 1. However, the curvatures are very closely distributed around the median, and a priori, this is not a very favorable situation to select a given link for surgery. The correlation of this graph is $38\%$, which means that there is an average edge-weight of 0.78. In Fig. 1 we can see the distributions of the edge weights and they are shifted downwards from 1 and are all very tightly distributed around a median value of 0.8, in accordance with the correlation. Here the $\kappa_{ij}$ are all positive which corresponds to a graph of positive curvature, while the edge weights are not identical in a fully connected graph, due to the connectivity. With the Ricci coarse curvature, we are able to use the full correlation matrix to build a network and carry our Ricci flow without any assumptions about the number of clusters.
@@
We expect that Ricci flow should shrink the edges of a community to regions of large positive curvature. This is indeed the case. If we use edge weights as a measure, it is still not clear how to carry out surgery in the fully connected manifold using distances. In [26] edge curvature alone without Ricci flow was used, with the argument that triangles are a good index of community formation. We suggest in this work, based on Ricci flow and the intuition obtained from the fully connected graph, that curvature is a good metric to use.
@@
From Lin and Yau[13],[37], the infinite weighted graph has a lower bound to the ORC given by $\kappa_{ij}\geq\kappa_{ij,min}=\frac{2}{d_{i}}+\frac{2}{d_{j}}-2$, where $d_{i}$ is the degree of link $i$. The lower bound has been further developed for hypergraphs[38]. In the limit of the infinite complete graph with equal edge weights, this gives a value for $\kappa_{ij,min}=\frac{2(3-n)}{(n-1)}$. For the NASDAQ-100 graph this gives $\kappa_{ij,min}\approx-1.96$. Naively, this would correspond to a length of $d_{ij}\approx 2.96$ as the longest edge length possible. We find after 20 iterations that the graph splits into two communities, with very short edge weights and positive curvature. These two communities are separated by long, flat links with curvature zero and edge weights around 3. It is interesting to note that the fully connected graph which satisfies strict positivity of the curvature before the Ricci flow, still obeys the lower bound condition, even with unequal edge weights. This brings us to the question of surgery in such a graph, it would seems appropriate to use the flattening that develops along the neck pinch singularity to perform surgery along the zero-curvature links. This suggests that the geometry of the NASDAQ-100 graph is a prominent influence on its properties. The reason is that the degree of the graph seems to play a key role in influencing the lower bound, even without equal edge weights.

### v2

@@ Abstract
We use the discrete Ollivier-Ricci graph curvature with Ricci flow to examine the intrinsic geometry of financial markets through the empirical correlation graph of the NASDAQ-100 index. Our main result is the development of a technique to perform surgery on the neckpinch singularities that form during the Ricci flow of the empirical graph, using the behavior and the lower bound of curvature of the fully connected graph as a starting point. We construct an algorithm that uses the curvature generated by intrinsic geometric flow of the graph to detect hidden hierarchies, community behavior, and clustering in financial markets despite the underlying challenges posed by a highly connected geometry.
@@
Edge-centric approaches that involve graph curvature are increasingly emerging as the method of choice to analyze complex graphs. Combinatorial Forman-Ricci[19],[14],[20] curvature and probabilistic Ollivier-Ricci curvature[8],[9],[10] (ORC) are key developments that have allowed the adaptation of differential geometry methods to discrete structures such as graphs. The correlation matrix, being a fully connected graph, already provides an important indicator of the geometry of the system. Graph curvature has been evoked in the context of the stock market, but in these works the network was again replaced by MST[21],[22],[23]. The continuum approach to ORC was analysed in [24]. In this letter, we use the Ollivier-Ricci curvature to construct and analyze the weighted network generated by the stock market as a weighted complete graph. Our main result is to propose a discrete Ricci flow algorithm with surgery based on curvature along neckpinch singularities, which allows us to analyze the structure and intrinsic geometry of fully connected graphs with positive curvature, such as the stock market. Detailed comparisons with state-of-the art techniques such as MST and PCA are the subject of a longer paper[25].
@@
From this definition, we see that if the nodes $i$ and $j$ belong to the same community, they tend to share neighbors. Therefore, the cost of transporting the mass is smaller and the Wasserstein distance is not greater than $d(i,j)$. Thus, intra-community edges have a positive Ollivier-Ricci curvature. On the other hand, if nodes $i$ and $j$ are from different communities, they tend to have fewer neighbors in common. Therefore, transporting the mass from $i$ to $j$ necessarily involves passing through the edge $ij$ and the Wasserstein distance is greater than $d(i,j)$. Therefore, the curvature of the edge $ij$ is negative and the link is hyperbolic. On a flat graph, the Wasserstein and $d(i,j)$ distances are identical and the curvature is zero. If all the links in a graph have zero curvature, then the graph is called Ricci-flat. The regular grid is an example of a Ricci-flat graph.
@@
In Fig. 1 we present our construction of the NASDAQ-100 graph with empirical edge weights, constructed from real-life data. We recovered the values of stock prices from Yahoo! Finance[32] using the components of the NASDAQ-100 Index from [33]. We calculate the returns and average over a period of five years from 1/1/2020-20/11/2024. Based on the edge weights, we obtain a histogram of Ricci curvatures[27],[34]. The histogram is shifted downward from the complete graph at 0.505, though being all positive retain the spherical aspect of this graph. However, the curvatures are very closely distributed around the median, and a priori, this is not a very favorable situation to select a given link for surgery. The correlation of this graph is $38\%$, which means that there is an average edge-weight of 1.11. In Fig. 1 we can see that the distributions of the edge weights are all tightly distributed around a median value of 1.11, in accordance with the correlation. Here the $\kappa_{ij}$ are all positive which corresponds to a graph of positive curvature, while the edge weights are not identical in a fully connected graph, due to the connectivity. With the Ricci coarse curvature, we are able to use the full correlation matrix to build a network and carry our Ricci flow without any assumptions about the number of clusters.
@@
We expect that the Ricci flow should shrink the edges of a community to regions of large positive curvature. This is indeed the case. If we use edge weights as a measure, it is still not clear how to carry out surgery in the fully connected manifold using distances. In [26] edge curvature alone without Ricci flow was used, with the argument that triangles are a good index of community formation. We suggest in this work, based on Ricci flow and the intuition obtained from the fully connected graph, that curvature is a good metric to use.
@@
The Ollivier-Ricci graph curvature was shown to be bounded by -2 and 1, with the bounds attained for the infinite double stars and complete graphs[13],[37]. Therefore, we see that the curvature is subject to strong constraints. From Lin and Yau[13],[37], the lower bound of the ORC is given by $\kappa_{ij}\geq\kappa_{ij,min}=\frac{2}{d_{i}}+\frac{2}{d_{j}}-2$, where $d_{i}$ is the degree of link $i$. The lower bound has been further developed for hypergraphs[38]. In the limit of equal edge weights, this gives a value for $\kappa_{ij,min}=\frac{2(3-n)}{(n-1)}$. For the NASDAQ-100 graph this gives $\kappa_{ij,min}\approx-1.96$. Naively, this would correspond to a length of $d_{ij}\approx 2.96$ for the edge length. For the positive case, the edge length shrinks as expected, while the edge length increases for the negative (hyperbolic case). The edge length will continue to increase in the hyperbolic case even when the theoretic lower bound of -2 is attained. Therefore identifying the lower bound of the curvature is a more reliable index for community detection. We find after 20 iterations that the graph splits into two communities, with very short edge weights and positive curvature. These two communities are separated by long, flat links with curvature zero and edge weights around 3. It is interesting to note that the fully connected graph which satisfies strict positivity of the curvature before the Ricci flow, still obeys the lower bound condition, even with unequal edge weights. This brings us to the question of surgery in such a graph, it would seems appropriate to use the flattening that develops along the neck pinch singularity to perform surgery along the zero-curvature links. This suggests that the geometry of the NASDAQ-100 graph is a prominent influence on its properties. The reason is that the degree of the graph seems to play a key role in influencing the lower bound, even without equal edge weights.